# Multi-Step Embed to Control: A Novel Deep Learning-based Approach for Surrogate Modelling in Reservoir Simulation

### v1

@@ 4.1 Comparison of E2C and Multi-step E2C
Finally, to demonstrate the effectiveness of the proposed multi-step E2C across various testing scenarios, Figure 7 presents a boxplot that shows the absolute error of 100 testing cases over time for different approaches. As the plot illustrates, the E2C approach consistently exhibits the highest absolute error across timesteps, indicating it is less accurate than the other methods. In contrast, the multi-step E2C approaches with k=5 and k=7 show similar accuracy, with the multi-step E2C (k=5) performing slightly better. This slight performance difference can be attributed to the reduction in training samples when the forward timesteps are increased, making the network more challenging to train effectively.

### v2

@@ 3 Methods
In this proposed approach, it is assumed that a sequence of states $x_{t}$, $x_{t+1}$, …, $x_{t+k}$ can be generated by corresponding latent states $z_{t}$, $z_{t+1}$, …, $z_{t+k}$ with the decoder structure. The key distinction of our method lies in how the latent space is propagated. [21] attempted to use the E2C approach but found that special treatment was needed to maintain stability. Unlike the locally linear form used in E2C, where locally A and B matrices are applied to transitions (by locally we meant that A and B matrices in 4 change with different latent states), our model leverages the Koopman operator to project the latent space into a higher-dimensional space where multi-step transitions become globally linear, as in 7. As a result, our model, referred to multi-step embed to control (MSE2C), models the likelihood $p(x_{t},x_{t+1},...x_{t+k}|u_{t},u_{t+1},...u_{t+k-1})$, where $k$ is the forward step. The resulting loss of the whole framework $\mathcal{L}_{total}$ is written as follows(the details are presented in next section):
$\mathcal{L}_{total}=\mathcal{L}_{rec}+\mathcal{L}_{reg}+\mathcal{L}_{koop}$ (5)
@@ 4.1 Comparison of E2C and Multi-step E2C
Finally, to demonstrate the effectiveness of the proposed multi-step E2C across various testing scenarios, Figure 7 presents a boxplot that shows the absolute error of 100 testing cases over time for different approaches. As the plot illustrates, the E2C approach consistently exhibits the highest absolute error across timesteps, indicating it is less accurate than the other methods. In contrast, the multi-step E2C approaches with k=5 and k=7 show similar accuracy, with the multi-step E2C (k=5) performing slightly better. This slight performance difference can be attributed to the reduction in training samples when the forward timesteps are increased(as presented in 8), making the network more challenging to train effectively.
@@
Figure 8: Collected samples decreases as forward step $k$ increases.